# Metric Graph Kernels via the Tropical Torelli Map

### v1

@@ 5.3 Urban Road Network Classification
Urban road networks (URNs) are naturally modeled as metric graphs, where a combinatorial model is built on top of landmarks (nodes) and roads (edges). These networks encode both geometric and topological information, capturing the structural organization of cities. Global URNs present dominant patterns across multiple scales, which are closely related to continents and regions. Based on the study of [10], we focus on URN patterns at small scale ($<$ 500 meters), grouped into 4 categories: gridiron, linear, chaotic, tributary. Figure 4 presents sample cities from each category.
Constructing Datasets. The spatial data of URNs can be retrieved from the OpenStreetMap (OSM) database [23], queried by the open source Python library OSMnx [3].
We constructed 8 datasets for classification in the following way: First, we picked a city for each category. For a city with latitude–longitude coordinate $C=(x,y)$, we retrieved a large road network centered at $C$ with global radius $R$. We then randomly sampled $L$ landmarks from the large road network. For each landmark, we retrieved a small road network with local radius $r$. Full details are given in Table 8 in Appendix D.1.
Classification Results. We performed classification using C-SVM for 4 graph kernels: OS, SP, TTE, and TTW. Figure 5 displays an exploratory dimension reduction analysis via kernel PCA, applied to the datasets URN2-1S and URN3-S with the TTE and TTW kernels.
Table 2 shows the complete list of dataset statistics, classification accuracy and computational time. The TTE and TTW kernels outperform the other graph kernels on most datasets. Notably, URNs are typically sparse, which makes the TTE and TTW kernels particularly well-suited for these datasets, especially for large graphs where other kernels fail due to memory limitations. Additional experimental results can be found in Appendix D.2.
(a) URN2-1S
(b) URN3-S
@@ 6 Discussion
We proposed new graph kernels which are extendable to the space of metric graphs. We adapted the classical Torelli map from tropical geometry to weighted graphs, embedding them into the space of PSD matrices. We constructed our kernels using two distance functions. Through comprehensive experiments, we demonstrated the effectiveness and applicability of our graph kernels.
Limitations and Future Work. Graphs that are trees have genus 0, so the tropical Torelli map is trivial. Consequently, the resulting kernels identify all trees. Also our method does not incorporate label or attribute information. Developing a principled way to define and integrate labels on metric graphs remains an open challenge and is a promising direction for future work.
Appendix A Further Details of the Tropical Torelli Map
A.1 Geometric Background
@@
In classical algebraic geometry over the complex numbers, one of the fundamental results is that every smooth projective algebraic curve $X$ of genus $g$ encodes rich geometric information in a special complex torus $\mathbb{C}^{g}/L$. This object is known as the Jacobian variety of the curve $\mathrm{Jac}(X)$. The classical Torelli map is a natural and powerful construction that assigns to each curve $X$ its Jacobian $\mathrm{Jac}(X)$. This map is injective, meaning that the Jacobian retains enough information to uniquely recover the curve. The Torelli map thus plays a central role in the study of algebraic curves by linking their intrinsic geometry to the rich structure of abelian varieties [17].
In tropical geometry—a piecewise-linear analog of algebraic geometry—a tropical curve is represented not by a smooth curve, but by a metric graph $\Gamma$. Just as classical algebraic curves have associated Jacobian varieties, tropical curves have the tropical Jacobian variety, which is a real torus of the form $\mathrm{Jac}(\Gamma)=\mathbb{R}^{g}/L$. The tropical Torelli map is the function that assigns to each tropical curve $\Gamma$ to its corresponding tropical Jacobian $\mathrm{Jac}(\Gamma)$. This map captures the essential geometric and topological information of the graph in a purely combinatorial and metric framework [6].
In computations we have to choose a vector space basis for $\mathbb{R}^{n}$ and a lattice basis for $L$, hence the tropical Jacobian can be represented in the following equivalent forms:
  • (I)
    The tropical form $(\mathbb{R}^{g}/Q,Q^{-1})$, where the lattice is spanned by the column vectors of $Q$, and the inner product on $\mathbb{R}^{g}$ is represented by $Q^{-1}$. Under this form, the squared distance function is given by
    $d^{2}_{(\mathrm{I})}([x],[y])=\min_{w\in\mathbb{Z}^{g}}\big{(}(x-y-Qw)^{\top}Q^{-1}(x-y-Qw)\big{)}.$
    Form (I) is particularly useful for some computations in tropical geometry such as the tropical Abel–Jacobi transform of metric graphs [7].
  • (II)
    The standard lattice form $(\mathbb{R}^{g}/\mathbb{Z}^{g},Q)$, where the lattice is $\mathbb{Z}^{n}$, and the inner product on $\mathbb{R}^{g}$ is represented by $Q$. Under this form, the squared distance function is given by
    $d^{2}_{(\mathrm{II})}([x],[y])=\min_{w\in\mathbb{Z}^{g}}\big{(}(x-y-w)^{\top}Q(x-y-w)\big{)}.$
    Our main algorithm in Section 3.2 of the tropical Torelli map for weighted graphs is based on form (II).
  • (III)
@@
    The standard inner product form $(\mathbb{R}^{g}/\sqrt{Q},\|\cdot\|_{2})$ where the lattice is spanned by the column vectors of $\sqrt{Q}$, and the inner product on $\mathbb{R}^{g}$ is the standard $L_{2}$ inner product. Under this form, the squared distance function is given by
    $d^{2}_{(\mathrm{III})}([x],[y])=\min_{w\in\mathbb{Z}^{g}}\big{\|}x-y-\sqrt{Q}w\big{\|}_{2}^{2}.$
    Form (III) is the motivation for us to consider the Bures–Wasserstein distance in the construction of tropical Torelli–Wasserstein graph kernel in Section 4.2.
A.2 Proof of Theorem 1
@@
We first show that if $G$ has a generic length function, then $G$ can be oriented in a canonical way through the following operations: (1) for any $e\in T$ with endpoints $u,v$, define $\omega(u)$ as the length of the other tree edge incident to $u$ (set $\omega(u)=0$ if no such edge exists). Then orient $e$ from low $\omega$ value to the high $\omega$ value; (2) for any $e\notin T$ with endpoints $u,v$, define $w(u)$ as the length of the unique tree edge incident to $u$ that lies on the path from $u$ to $v$ in $T$. Then orient $e$ from low $\omega$ value to the high $\omega$ value.
@@
In Algorithm 1, the edges not in the minimal spanning $T$ are sorted by length. With a generic length function $\ell$, the minimal spanning tree $T$ is unique and the edge lengths are distinct, thus the 1-cycles $\sigma_{1},\ldots,\sigma_{g}$ are uniquely determined. It suffices to show that the cycle-edge matrix is independent of the ordering of edges in $T$. In fact, suppose $e^{\prime}_{1},\ldots,e^{\prime}_{n-1}$ is another ordering of edges in $T$. Then there exists a permutation matrix $P\in S_{n-1}$ such that $M_{T}^{\prime}=M_{T}P$ and $L_{T}^{\prime}=P^{\top}L_{T}P$. Therefore,
@@
Then we show that the matrix $Q$ is invariant under edge subdivision. Without loss of generality we assume that $e_{s}$ is subdivided into $e^{\prime}_{s}$ and $e_{s+1}^{\prime}$. For remaining edges we set $e^{\prime}_{k}=e_{k}$ if $k<s$, and $e^{\prime}_{k}=e_{k-1}$ if $k>s+1$. For cycle-edge incidence matrices we have
[…]

### v2

@@ 5.3 Urban Road Network Classification
Urban road networks (URNs) are naturally modeled as metric graphs, where a combinatorial model, i.e., a weighted graph, is built on top of landmarks (nodes) and roads (edges). These networks encode both geometric and topological information, capturing the structural organization of cities. Global URNs present dominant patterns across multiple scales, which are closely related to continents and regions. Based on the study of Chen et al. (2024), we focus on URN patterns at small scale (<< 500 meters), grouped into 4 categories: gridiron, linear, chaotic, tributary. Figure 3 presents sample cities from each category.
@@
Constructing Datasets. The spatial data of URNs were retrieved from the OpenStreetMap (OSM) database (OpenStreetMap contributors, 2017), queried by the open source Python library OSMnx (Boeing, 2025). We constructed 8 datasets for classification: First, we picked a city for each category. For a city with latitude–longitude coordinate $C=(x,y)$, we retrieved a large road network centered at $C$ with global radius $R$ and then randomly sampled $L$ landmarks from the large network. For each landmark, we retrieved a small road network with local radius $r$. See Table 8 in Appendix F.1 for full details.
(a) URN2-1S
(b) URN3-S
Classification Results. We performed classification using C-SVM for 4 graph kernels: OS, SP, TTE, and TTW. Figure 4 displays an exploratory dimension reduction analysis via kernel PCA. Table 2 shows the complete list of dataset statistics, classification accuracy and computational time. The TTE and TTW kernels outperform the other graph kernels on most datasets. Notably, URNs are typically sparse, which makes the TTE and TTW kernels particularly well-suited for these datasets, where other kernels fail due to memory limitations. Additional results can be found in Appendix F.2.
@@ 6 Discussion
We proposed new graph kernels which are extendable to the space of metric graphs. We adapted the classical Torelli map from tropical geometry to weighted graphs, embedding them into the space of PSD matrices. We constructed our kernels using two distance functions. Through comprehensive experiments, we demonstrated the effectiveness and applicability of our metric graph kernels.
Limitations and Future Research. Graphs that are trees have genus 0, so the tropical Torelli map is trivial. Consequently, the resulting kernels identify all trees. There are several possible directions to modify either the graph or the graph kernel to incorporate tree-like information, which we discuss further in Appendix G.
Our method does not incorporate label or attribute information. However, as discussed in Section 2, the correct analogue of labels in the setting of metric graphs should be functions defined on the underlying metric space. Developing a principled way to define and integrate such definition on metric graphs remains an open challenge.
Finally, we emphasize that there currently exist no standard nor appropriate benchmark datasets for evaluating methods for metric graphs. Many real-world graphs are not naturally endowed with a meaningful metric structure: for example, molecular graphs encode chemical bonds between atoms, but these bonds do not correspond to geometric edges with interpretable lengths. Tested datasets in this work, such as AIDS, BZR, and COX2 fall into this category. Similarly, edge weights or distances in social networks lack a coherent geometric interpretation. Building principled metric graph benchmark datasets is thus an important line for future work.
Acknowledgments
Y.C. is supported by Digital Futures Postdoctoral Fellowship. A.M. is supported by the EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Background on the Tropical Torelli Map
In this section of the Appendix, we provide further background on the tropical Torelli map. We present its original definition in both the coordinate-free version and matrix versions. We also provide further discussions on the tropical Torelli map for weighted graphs, especially its motivations.
A.1 The Coordinate-Free Version
@@
In classical algebraic geometry over the complex numbers, one of the fundamental results is that every smooth projective algebraic curve $X$ of genus $g$ encodes rich geometric information in a special complex torus. This object is known as the Jacobian variety of the curve $\mathrm{Jac}(X)$. The classical Torelli map is a natural and powerful construction that assigns to each curve $X$ its Jacobian $\mathrm{Jac}(X)$. To study this map, appropriate moduli spaces for its domain and codomain are required. In classical algebraic geometry, the construction and study of such moduli spaces has been a central theme of research. More precisely, the Torelli map is defined as a morphism $\mathcal{M}_{g}\to\mathcal{A}_{g}$, where $\mathcal{M}_{g}$ is the moduli space of smooth complex projective curves of genus $g$, and $\mathcal{A}_{g}$ is the moduli space of principally polarized abelian varieties of dimension $g$. A fundamental result, known as the Torelli theorem, asserts that this map is injective, meaning that the Jacobian retains enough information to uniquely recover a smooth projective curve. The Torelli map thus plays a central role in the study of algebraic curves by linking their intrinsic geometry to the rich structure of abelian varieties (Miranda, 1995).
In tropical geometry, an (abstract) topical curve $\Gamma$ is defined as a metric graph. It can be obtained from a particular degeneration process of a one-parameter family of smooth curves, where the edge lengths measure the speed of such degeneration (Chan, 2021). Analagous to the classical setting, every tropical curve is canonically associated with a real torus called its tropical Jacobian $\mathrm{Jac}(\Gamma)$. Using graph homology, overviewed in Section 3.1, the tropical Jacobian is defined as the quotient $H_{1}(\Gamma;\mathbb{R})/H_{1}(\Gamma;\mathbb{Z})$, equipped with a canonical positive definite quadratic form induced by the edge lengths of $\Gamma$, which plays the role of a tropical principal polarization.
The tropical Torelli map assigns to each tropical curve $\Gamma$ its corresponding tropical Jacobian $\mathrm{Jac}(\Gamma)$. As in the classical case, a meaningful study of this map requires the introduction of appropriate moduli spaces for its domain and codomain. In contrast to the classical construction over the complex numbers, which relies heavily on complex geometry, the moduli space $\mathcal{M}_{g}^{\mathrm{trop}}$ of tropical curves of genus $g$ and the moduli space $\mathcal{A}_{g}^{\mathrm{trop}}$ of principally polarized tropical abelian varieties of dimension $g$ are constructed using combinatorial and polyhedral methods. With these constructions, the tropical Torelli map is defined as a morphism $\mathcal{M}_{g}^{\mathrm{trop}}\to\mathcal{A}_{g}^{\mathrm{trop}}$. It can be shown that the tropical Torelli map is injective on the locus of 3-edge-connected tropical curves (Brannetti et al., 2011).
A.2 The Matrix Version
Notice that, in the coordinate-free definition, the tropical Torelli map assigns to a tropical curve a torus, which is not well suited for explicit computation since it is an abstract geometric object. For the purpose of computation and using this construction in applications, as is our goal in this paper, it is therefore necessary to fix a concrete representation of the tropical Jacobian. At the same time, we must be mindful of the situation when different representations correspond to isomorphic tropical Jacobians, and hence determine the same point in the target of the tropical Torelli map.
@@ A.2 The Matrix Version
For any tropical Jacobian, there exist bases under which it can be represented in the standard form $(\mathbb{R}^{g}/\sqrt{Q},\|\cdot\|_{2})$, where the lattice is spanned by the column vectors of $\sqrt{Q}$, and the tropical polarization is represented by the standard Euclidean inner product. Under this form, the squared distance function is given by
$d^{2}([x],[y])=\min_{w\in\mathbb{Z}^{g}}\big\|x-y-\sqrt{Q}w\big\|_{2}^{2}.$
Another torus $(\mathbb{R}^{g}/\sqrt{Q^{\prime}},\|\cdot\|_{2})$ is isomorphic to $(\mathbb{R}^{g}/\sqrt{Q},\|\cdot\|_{2})$ if and only if there exists an invertible integer matrix $A\in\mathrm{GL}_{g}(\mathbb{Z})$ and an orthogonal matrix $U\in O(g)$ such that $\sqrt{Q^{\prime}}=U\sqrt{Q}A$. This representation of the tropical Jacobian is the motivation for many constructions of our metric graph kernel.
Given this observation, we are now able to define the tropical Torelli map by sending each tropical curve $\Gamma$ to the equivalence class $Q(\Gamma)]$, where the equivalence relation is defined by the isomorphism described above. This definition could be further simplified if we can always select a unique representative from the class $Q(\Gamma)]$, however, such a choice is not available in general without imposing additional assumptions.
A.3 The Tropical Torelli Map for Weighted Graphs
From a computational perspective, there are two important restrictions arising from the original definition of the tropical Torelli map:
  • The map is defined on the moduli space of tropical curves of fixed genus $g$. Although the setting is natural for theoretical considerations, it is not well suited if we want to compare metric graphs of different genus.
  • Even in the matrix version of the tropical Torelli map, working with entire equivalence classes of matrices is computationally prohibitive, as it involves the arithmetic group $\mathrm{GL}_{g}(\mathbb{Z})$, which is discrete and infinite, and therefore standard continuous optimization methods cannot be applied.
The first restriction can be disregarded if the goal is simply to compute the image of the tropical Torelli map for a given graph. To address the second restriction, a practical approach is to fix a basis a priori, which is equivalent to choosing a spanning forest of a weighted graph (or a spanning tree if the graph is connected) and ordering the remaining non-tree edges. With such a choice given, the tropical Torelli map can be well defined for a weighted graph $G$ with a distinguished spanning forest $\mathrm{SF}_{G}$.
Definition A.1.
Let $\mathcal{M}^{*}_{\leq g}$ be the set of weighted graphs of genus at most $g$, equipped with distinguished spanning forests. The tropical Torelli map for weighted graphs is given by $\mathcal{T}:\mathcal{M}^{*}_{\leq g}\to\mathrm{SPD}(g),\,G\mapsto Q(G)$, where $Q(G)$ is determined by the spanning forest $\mathrm{SF}_{G}$.
A generic length function on a weighted graph canonically determines a spanning forest, namely the minimal one (Definition 3.1). In this sense, the condition of a generic length function on a weighted graph is stronger and implies the condition of a distinguished spanning forest.
Since the bases of the tropical Jacobian are fixed a priori, there is no need to consider equivalence under the arithmetic group $\mathrm{GL}_{g}(\mathbb{Z})$, which significantly simplifies the computation. Notice that without specifying the bases, a possible well-defined distance between $Q_{1}]$ and $Q_{2}]$ would be
$d([Q_{1}],[Q_{2}])=\min_{\begin{subarray}{c}U_{1},U_{2}\in O(g)\\ A_{1},A_{2}\in\mathrm{GL}_{g}(\mathbb{Z})\end{subarray}}\|U_{1}\sqrt{Q_{1}}A_{1}-U_{2}\sqrt{Q_{2}}A_{2}\|,$
for which it is difficult to obtain a closed-form expression or to minimize numerically. However, in the absence of $\mathrm{GL}_{g}(\mathbb{Z})$, the resulting distance is exactly the Bures–Wasserstein distance (Section 4.2).
A.4 Example
(a) Example graph
(b) Orientation and minimal spanning tree
Figure 5: Computation of the tropical Torelli map for a weighted graph. On the right panel, the orientation is indicated by arrow and the minimal spanning tree is colored in red.
We demonstrate the computation of the tropical Torelli map for weighted graphs on a toy example shown in Figure 5(a). The graph $G$ has 6 nodes and 8 edges. The length of each edge is defined as $\ell(e_{i})=i$. We compute the minimal spanning tree as $T=\{e_{1},e_{2},e_{4},e_{5},e_{6}\}$, which is shown in Figure 5(b). After fixing an orientation, each edge not in $T$ determines a 1-cycle as follows:
$\sigma_{1}$ $=e_{3}-e_{2}+e_{1},$
$\sigma_{2}$ $=e_{7}-e_{5}-e_{6},$
$\sigma_{3}$ $=e_{8}-e_{6}-e_{4}-e_{2}.$
Thus the cycle-edge incidence matrix is given by
The matrix of the tropical Torelli map can be computed by
$Q=MLM^{\top}=\begin{bmatrix}6&0&2\\ 0&18&6\\ 2&6&20\end{bmatrix}$.
Appendix B Background on Kernel Support Vector Machines
In this section, we review the fundamentals of kernel Support Vector Machines (SVMs), which are used in our graph classification experiments. We also provide background on kernel functions, in particular on their construction from distance functions and their positive definiteness.
B.1 A Short Review on Support Vector Machines
General Setup. Support Vector Machines (SVMs) are supervised learning methods for classification and regression that construct decision functions based on margin maximization (Steinwart and Christmann, 2008). In the binary classification setting, we are given training data $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ where $x_{i}\in\mathbb{R}^{m}$ and $y_{i}\in\{+1,-1\}$, the central idea of SVM is to find a hyperplane $f(x)=w^{\top}x+b$ that separates the two classes with the maximum margin, i.e., the maximum distance between the hyperplane and the closest data points. For linearly separable data, we have the so-called hard-margin maximization, which can be formulated as
$\min_{w,b}\frac{1}{2}\|w\|^{2}\quad\text{ subject to}\quad y_{i}(w^{\top}x_{i}+b)\geq 1,\forall i\in\{1,\cdots,n\}.$
Soft-Margin Formulation. In practice, exact separability rarely holds. This motivates the soft-margin formulation, which is commonly known as C-SVM. By introducing slack variables $\xi_{i}\geq 0,i=1,\ldots,n$, the C-SVM optimization problem is
[…]
which is equivalent to the following unconstrained problem with hinge loss:
$\min_{w,b}\frac{1}{2}\|w\|^{2}+C\bigg[\sum_{i=1}^{n}\max(0,1-y_{i}(w^{\top}x_{i}+b))\bigg].$
In the case of C-SVM, the hyperparameter $C>0$ is a regularization parameter controlling the trade-off between margin maximization and constraint violations. Large $C$ adds more penalty on misclassification while small $C$ yields wider margins at the expense of allowing more violations.
Prediction. The Lagrangian dual of (2) is given by
[…]
The optimal multipliers $\alpha^{*}$ can be obtained by solving (3), and the primal-dual relationship gives
$w=\sum_{i=1}^{n}\alpha^{*}_{i}y_{i}x_{i}.$
Only those samples with $\alpha^{*}_{i}>0$ contributes to $w$, which are called support vectors. The bias term can be obtained using any support vector by setting $b=y_{i}-\sum_{j=1}^{n}\alpha_{j}^{*}y_{j}x^{\top}_{j}x_{i}$. Then the decision function is given by
$f(x)=w^{\top}x+b=\sum_{i=1}^{n}\alpha^{*}_{i}y_{i}x_{i}^{\top}x+b.$
For any new data $x_{\mathrm{new}}$, the prediction is $\hat{y}=\mathrm{sign}(f(x_{\mathrm{new}}))$.
The Kernel Trick. The dual formulation (3) only depends on the inner product $x_{i}^{\top}x_{j}$. This observation leads to the so-called kernel trick which essentially replaces $x_{i}^{\top}x_{j}$ by a kernel function $k(x_{i},x_{j})$. Therefore the kernel SVM formulation is given by
[…]
The decision function then becomes
$f(x)=\sum_{i=1}^{n}\alpha^{*}_{i}y_{i}k(x_{i},x)+b,$
with the optimal multipliers $\alpha^{*}$ solved from (4).
B.2 Background on Kernels
Kernels form a foundational class of similarity functions in machine learning and data analysis, which are often used to capture nonlinear relationships by embedding data into high-dimensional feature spaces. Let $\mathcal{X}$ be a set and $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a symmetric function. Commonly $k$ is required to be positive definite in the sense that for any finite set of points $x_{1},\ldots,x_{n}\in\mathcal{X}$ the matrix $K_{ij}=k(x_{i},x_{j})]_{n\times n}$ is positive semi-definite, i.e., for any real numbers $c_{1},\ldots,c_{n}$, we have
$\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}k(x_{i},x_{j})\geq 0.$
In the kernel SVM formulation, a positive semi-definite matrix $K_{ij}]_{n\times n}$ ensures that (4) is a quadratic program.
For any positive definite kernel $k$, there exists a unique Hilbert space $\mathcal{H}$ known as the reproducing kernel Hilbert space (RKHS), along with a map $\phi:\mathcal{X}\to\mathcal{H}$ called the feature map, such that the value of the kernel is given by the inner product of the feature map (Wainwright, 2019):
$k(x,y)=\langle\phi(x),\phi(y)\rangle_{\mathcal{H}}.$
A typical way to construct a kernel involves the distance on the underlying space. Let $d(x,y)$ be a distance function on $\mathcal{X}$, the distance is said to be of conditionally negative type if for any finite set of points $x_{1},\ldots,x_{n}\in\mathcal{X}$ and any real numbers $c_{1},\ldots,c_{n}$ such that $\sum_{i=1}^{n}c_{i}=0$, we have
$\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}d(x_{i},x_{j})\leq 0.$
If $d(x,y)$ is of conditionally negative type, then the kernel defined by $k(x,y)=\exp(-\gamma d(x,y)^{2})$ is positive definite for any $\gamma>0$; this result is known as Schoenberg’s theorem (Sejdinovic et al., 2013). The Euclidean distance $d(x,y)=\|x-y\|_{2}$ for $x,y\in\mathbb{R}^{m}$ is a typical example of a distance of conditionally negative type. However, the 2-Wasserstein distance is not of conditionally negative type on the space of Gaussians of dimension greater than 1 (Peyré et al., 2019), and thus the induced kernel is indefinite. In practice, indefinite kernels are also widely used, and the corresponding learning methodology is called indefinite learning. In general, an indefinite kernel corresponds to a Reproducing Kernel Kreĭn Space (RKKS), on which many RKHS-based techniques can be extended (Oglic and Gärtner, 2018; Togninalli et al., 2019).
Appendix C Technical Proofs
C.1 Proof of Theorem 3.2
Proof.
@@ Appendix C Technical Proofs
We first show that if $G$ has a generic length function, then $G$ can be oriented in a canonical way through the following operations: (1) for any $e\in T$ with endpoints $u,v$, define $\omega(u)$ to be the length of the other tree edge incident to $u$. We set $\omega(u)=0$ if no such edge exists; if there is more than one incident edge, we choose $\omega(u)$ to be the minimum length among all incident tree edges. Then we orient $e$ from low to high $\omega$ value; (2) for any $e\notin T$ with endpoints $u,v$, define $w(u)$ as the length of the unique tree edge incident to $u$ that lies on the path from $u$ to $v$ in $T$. Then we orient $e$ from low to high $\omega$ value.
@@
In Algorithm 1, the edges not in the minimal spanning $T$ are sorted by length. With a generic length function $\ell$, the minimal spanning tree $T$ is unique and the edge lengths are distinct, thus the 1-cycles $\sigma_{1},\ldots,\sigma_{g}$ are uniquely determined. It suffices to show that the cycle-edge matrix is independent of the ordering of edges in $T$. Suppose $e^{\prime}_{1},\ldots,e^{\prime}_{n-1}$ is another ordering of edges in $T$. Then there exists a permutation matrix $P\in S_{n-1}$ such that $M_{T}^{\prime}=M_{T}P$ and $L_{T}^{\prime}=P^{\top}L_{T}P$, and
@@
We now show that the matrix $Q$ is invariant under edge subdivision. Without loss of generality we assume that $e_{s}$ is subdivided into $e^{\prime}_{s}$ and $e_{s+1}^{\prime}$. For the remaining edges, we set $e^{\prime}_{k}=e_{k}$ if $k<s$, and $e^{\prime}_{k}=e_{k-1}$ if $k>s+1$. For cycle-edge incidence matrices we have
$\left\{\begin{array}[]{ll}M^{\prime}_{ik}=M_{ik},&1\leq i\leq g\,,1\leq k\leq s,\\ M^{\prime}_{ik}=M_{i(k-1)},&1\leq i\leq g,\,s+2\leq k\leq m+1,\\ M^{\prime}_{i(s+1)}=M_{is},&1\leq i\leq g.\end{array}\right$.